part A .

Use a table of Standard Reduction Potentials to predict if a reaction will occur between Hg metal and Cl2(g), when the two are brought in contact via standard half-cells in a voltaic cell.

If a reaction will occur, write a balanced net ionic equation for the reaction, assuming that the product ions are in aqueous solution.

part B .

Enter electrons as e-.
Use smallest possible integer coefficients.
If a box is not needed, leave it blank.



Use the table 'Standard Reduction Potentials' located in the 'Tables', to predict if a reaction will occur between Hg metal and Cl2(g), when the two are brought in contact via half-cells in a voltaic cell.

If a reaction will occur, write a balanced net ionic equation for the reaction
Answer #1

E0 (Cl2 /Cl-) = + 1.36 V

and

E0 (Hg^2+ / Hg) = 0.78 V

as the reduction potential of Hg2+ / Hg is less than Cl2 / Cl-.

so

reaction occurs.

E0cell = 1.36 - 0.78 = 0.58 volt

as E0 cell is positive and hence reaction occurs.

Balanced net ionic equation is

Hg(l) + Cl2 (g) .................> Hg^2+ (aq) + 2 Cl- (aq)

actually HgCl2 should be precipitated as HgCl2 (s) but as mention above product ions in aqueous solution and so written as ions in aqueous phase.